Session: May 2001 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (A. Cattani at Thu Jun 7 02:06:22 METDST 2001) Minor failures

Torun (K.Borkowski at Thu Jun 7 09:32:06 METDST 2001) Minor failures
BBC5 and BBC8 continue malbehaviour (low levels; probably no useful data). No other known problems.

Effelsberg (A. Kraus at Thu Jun 7 09:53:14 METDST 2001) Success

Jodrell_Bank (Alastair Gunn at Thu Jun 7 11:26:39 METDST 2001) Success
No known problems.

Onsala (John Conway at Thu Jun 7 11:50:08 METDST 2001) Minor failures
Some passes show high parity errors, due posibility to vacumn column varaiations and shift of tape. Investigating.

Noto (P. Leto at Fri Jun 8 14:10:31 METDST 2001) Success

Westerbork (Tony Foley at Mon Jun 11 10:51:58 METDST 2001) Minor failures
No known errors